OpenStudy (anonymous):

Please help! Medal will be given. A parabola contains the points (1,8), (4, 29), and (-3, 8). What is the equation of this parabola in standard form?

OpenStudy (welshfella):

standard form is ax^2 + bx + c = 0 plug in the given values of x and y ( these are in the coordinates) to get 3 equations in a, b and c solve this sytem of equations and you have it

OpenStudy (anonymous):

@welshfella - I did that but I don't think I did it correctly. Here's my work: 8 = a + b + c 29 = 16a + 4b + c 8 = 9a - 4b + c ----------------------- 8 = a + b + c -c + 8 = a + b -c = -8 + a + b c = 8 - a - b ------------------------ 29 = 16a + 4b + 8 - a - b 29 = 15a +3b + 8 21 = 15a + 3b ------------------------ 8 = 9a - 4b + 8 - a - b 8 = 8a - 3b + 8 0 = 8a - 3b ------------------------- 21 = 15a + 3b -3b = -21 + 15a b = 7 + 15a (I made an error. It should be b = 7 - 5a) ------------------------- a = 0 b = 7 c = 1

OpenStudy (irishboy123):

8 = 9a - 4b + c ?????

OpenStudy (anonymous):

@IrishBoy123 - I think I did the problem wrong. Can you help me?

OpenStudy (irishboy123):

go back and correct that mistake (in line 3), the just follow ot through again, and you might well have got it right.

OpenStudy (anonymous):

@IrishBoy123 - It should be 8 = 9a - 3b + c, right?

OpenStudy (irishboy123):

yes.

OpenStudy (anonymous):

your 3rd equation is mistyped and it should be -3b instead of -4b: so you get a=1, b=2, c=5 then \[x ^{2}+2x+5\]

OpenStudy (anonymous):

@mlearning24 - Thank you.